OpenStudy (anonymous):

5(8 + 7) = 5 • x + 5 • 7

OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7

OpenStudy (anonymous):

75 + = 5x + 35 5x + 35 = 75 5x = 75-35 5x = 40 x = 40/5 = 8

OpenStudy (anonymous):

am I right

OpenStudy (anonymous):

Yes you're right.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Yw, good job on the explanation :)